Imagine rolling two fair 6 sided dice. What is the probability the sum of the rolls is 10?
Answer #1

Solution:

There are a total of 36 possible outcomes when we roll two fair 6 sided dice as shown in the figure below:

So,

There are three cases where the sum of the roles is exactly equal to 10, i.e. (6,4), (5,5) and (4,6)..

So in this case,

Probability = 3/36 = 1/12 = 0.0833...

End of the Solution...